Mum is 3 times as old as I am. she was 30 when I was born. what are our ages?
Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

m = mum's age.
s = son's age.
mum is 3 times the age of son.
equation for that is:
m = 3s
mum was 30 when son was born.
equation for that is:
m - s = 30
solve for m in this equation to get:
m = s + 30
in the equation of m = 3s, replace m with s + 30 to get:
s + 30 = 3s
subtract s from both sides of this equation to get:
30 = 2s
solve for s to get s = 30/2 = 15
since m = 3s, then m must be 45.
that should be y our answers.
mum is 45 years old and son is 15 years.
45 / 15 = 3 means mum is 3 times as old as son today.
45 - 15 = 0 means mum was 30 years old when she had son.

Answer by MathTherapy(10810)


Mum is 3 times as old as I am. she was 30 when I was born. what are our ages?
Let my age be I
Then mum's is: 3I
Since mum was 30 when I was born, mum will ALWAYS be 30 years older than me.
Therefore, we get: 3I = I + 30
Solve that for I to get my age. Add 30 to I to get mum's age!
That's it!!